Find the probability of the indicated event if P(E)=0.35 and P(F)=0.45

find P(E or F) if P(E and F) = 0.05

P(E or F)=


1
Expert's answer
2021-11-08T20:06:59-0500

"P(E or F)=P(E)+P(F)-P(E and F)=0.35+0.45-0.05=0.75."